OpenStudy (tennistar):

What is the exact value of the expression the square root of 16 times y to the sixth power. + 9the square root of y to the eighth power. - 3the square root of y to the sixth power. - the square root of y to the eighth power.

OpenStudy (tennistar):

\[\sqrt{16y ^{6}}+9\sqrt{y ^{8}}-3\sqrt{y ^{6}}-\sqrt{y ^{8}}\]

OpenStudy (anonymous):

\[\large \sqrt{16y^6}+9\sqrt{y^8}-3\sqrt{y^6}-\sqrt{y^8} \] like this?

OpenStudy (tennistar):

yes

OpenStudy (anonymous):

best thing to do is simplify it term by term... let's do the fist term: \(\large \sqrt{16y^6}=\sqrt{4^2(y^3)^2}=\sqrt{(4y^3)^2}=4y^3 \) can you do the others?

OpenStudy (tennistar):

No I am really stuck

OpenStudy (anonymous):

 ok... at least do B and C... use the pencil icon to continue drawing....